the supply and demand for broccoli are described by the following equations;supply QS=6p-80 Demand QD= 120-4p
a. grapgh the suppply curve and the demand curve .what is the equalibrium price and quantity?
b.calculate consumer surplus,producer surplus and total surplus at the equalibrium.
c.if a dictator who hated brocooli were to ban the vegetables,who would bear the larger burden_the buyers or sellers of brocooli?
1
Expert's answer
2020-05-08T14:54:42-0400

"a)6p-80=120-4p;"

"10p=200;"

"p=\$20;"

"Qs=6\u00d720-80=40;"

"Qd=120-4\u00d720=40;"


b)"Qd=120-4p=0;"

"4p=120;\np=30;"




  consumer surplus equal to the area of the hatched triangle.

consumer surplus"=\frac {1}{2}(30-20)*40=\$200;"


"Qs=6p-80=0;
p=\$13.3;"



producer  surplus equal to the area of the hatched triangle.

producer  surplus "=\frac{1}{2}(20-13.3)*40=" "\$134;"

"TS=CS+PS;
TS=134+200=\$334;"


c) If broccoli is banned, broccoli buyers will bear a greater burden.